# Kinematic cosmic dipole from a large sample of strong lenses

### 11 Mar 2026 v1

@@ 3.3 Lensing only results
While the dipole is detected in both the “100k” and “400k” samples for the highest peculiar velocity, $v_{o,\rm true}=1109.46$$\mathrm{km}\,\mathrm{s}^{-1}$}, the lowest CMB-inferred value does not produce a sufficiently strong signal to be detected in either sample. Fixing the dipole direction leads to a moderate tightening of the constraints on the dipole amplitude, improving the 68% credible interval width on $v_{o}$ by approximately 13–16%, depending on the sample size and fiducial velocity.
We therefore find that, when relying on lensing information alone, the Euclid lens sample is unlikely to provide strong discriminatory power between the CMB-inferred dipole amplitude and the larger values suggested by source number count analyzes. The resulting constraints are not sufficiently precise to exclude the CMB-inferred value beyond the $\sim 1.1\sigma$ level, even when fixing the dipole direction and adopting the most optimistic 400 000400\,000-lens scenario. We explore the possibility of including the stellar kinematic information to improve the precision of this measurement in the next section.
@@
Lensing only 100k $1608^{+1725}_{-1129}$ $263.6^{+64.9}_{-206.3}$ $10.6^{+44.2}_{-45.9}$ $1850^{+1741}_{-1254}$ $273.0^{+55.6}_{-202.9}$ $20.7^{+42.1}_{-39.8}$
Lensing only 400k $664^{+775}_{-465}$ $226.6^{+96.5}_{-175.7}$ $1.4^{+52.3}_{-49.8}$ $844^{+908}_{-597}$ $249.8^{+72.7}_{-180.3}$ $23.8^{+41.3}_{-45.7}$
Lensing only 100k - Fixed direction $1618^{+1367}_{-1035}$ $2111^{+1376}_{-1251}$
Lensing only 400k - Fixed direction $599^{+610}_{-410}$ $1023^{+751}_{-622}$
@@ 4.1 Spectroscopic velocity dispersion
$\theta^{\prime}_{E,\rm spec}=\frac{4\pi\sigma_{v}^{2}}{c^{2}}\frac{d[z_{l}^{\prime},z_{s}^{\prime}]}{d[z_{s}^{\prime}]}\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}+\frac{D_{LS}}{d[z_{l}^{\prime},z_{s}^{\prime}]}\frac{v_{o}}{c}-\frac{D_{S}}{d[z_{s}^{\prime}]}\frac{v_{o}}{c}\right),$ (4.11)
@@ 4.2 Fundamental Plane
$\log R_{e}$ $=a\log\sigma_{\mathrm{FP}}+b\log I_{e}+c\,.$ (4.13)
Only the angular size of the effective radius can be directly measured from the imaging data, but $R_{e}$ can be straightforwardly derived from this measurement and from the (photometric) redshift of the lens, assuming a cosmological model. We assume $a=1.177$, b=0.793b=-0.793 and $c=0.5$, obtained from the latest DESI measurement [Said2025]. Measurement errors on these parameters are ignored, as they lead to error much smaller than the intrinsic scatter of the FP. While $I_{e}$ is invariant under a boost (as shown in the App. D), $R_{e}$ is not. It can be computed from the observed subtended angle $\alpha^{\prime}$ and lens redshift $z_{l}^{\prime}$, which are affected by the boost, as
$R_{e}=\alpha d_{l}=\alpha^{\prime}d[z_{l}^{\prime}]\left(1+\frac{v_{o}}{2c}\cos(\theta_{\rm cm}^{\prime})+\frac{D_{L}}{d[z_{l}^{\prime}]}\frac{v_{o}}{c}\right)\,.$ (4.14)
@@
$\theta^{\prime}_{E,\rm FP}=\frac{4\pi\sigma_{\rm FP}^{\prime 2}}{c^{2}}\frac{d[z_{l}^{\prime},z_{s}^{\prime}]}{d[z_{s}^{\prime}]}\left(1-\frac{1}{2}\frac{v_{o}}{c}\cos(\theta^{\prime}_{\rm cm})+\frac{D_{LS}}{d[z_{l}^{\prime},z_{s}^{\prime}]}\frac{v_{o}}{c}-\frac{D_{S}}{d[z_{s}^{\prime}]}\frac{v_{o}}{c}+2\Sigma_{v}\frac{v_{o}}{c}\right)\,$, (4.15)
@@
$\Sigma_{v}$ $=\left(\frac{1}{2a}\cos(\theta_{\rm cm}^{\prime})+\frac{D_{L}}{ad[z_{l}^{\prime}]}\right)\,.$ (4.17)
@@
Details on how the uncertainties $\delta\sigma_{\mathrm{FP}}$ are calculated can be found in App. C. Similarly to Eq. (4.12), we introduced a bias and systematic error term to account for model mismatch, peculiar velocity effects, and redshift measurement errors when converting Einstein angles to velocity dispersions. We treat separately the likelihood contributions from lenses with only photometric redshift estimates and those with spectroscopic redshift measurements for both source and lens, since these two subsamples require different bias and systematic error corrections owing to the differing precision of their redshift determinations.
@@ 4.3 Euclid Forecast
Finally, we assess the constraining power of several spectroscopic follow-up scenarios for the Euclid strong-lens sample. In a first, pessimistic scenario, we assume that only 15 00015\,000 lenses receive spectroscopic follow-up, either from large spectroscopic surveys such as DESI [DESI2016] or 4MOST [4MOSToverview, 4SLSLS], or from dedicated follow-up programs. While both redshift and stellar velocity-dispersion measurements are in principle accessible with fiber spectroscopy, the latter are significantly more challenging, as they require a higher signal-to-noise ratio to robustly detect the absorption features of the lens galaxy. As a result, even if a lens is targeted by DESI or 4MOST, a substantial fraction of systems are expected to yield redshift measurements only. In this pessimistic scenario, we therefore assume that velocity-dispersion measurements are successfully obtained for only a third of the spectroscopically targeted lenses, i.e. 5 0005\,000 lenses. The remaining lenses are assumed to have spectroscopic, photometric, or no redshift information, as summarized in Table 3.
@@
Lensing only 100k $1436^{+1638}_{-1011}\ (0.5\sigma)$ $228.4^{+69.1}_{-115.4}$ $0.3^{+49.0}_{-48.6}$ $1629^{+1718}_{-1148}\ (0.7\sigma)$ $233.6^{+62.5}_{-108.6}$ $11.2^{+43.9}_{-46.7}$
FP (Photo-z) 100k $1090^{+959}_{-758}\ (1.1\sigma)$ $243.7^{+36.9}_{-46.9}$ $2.5^{+28.1}_{-27.3}$ $1713^{+895}_{-971}\ (1.0\sigma)$ $259.5^{+26.4}_{-29.3}$ $7.4^{+19.7}_{-16.7}$
FP (Spectro-z) 100k $463^{+521}_{-327}\ (1.2\sigma)$ $196.2^{+93.1}_{-111.1}$ $19.2^{+51.1}_{-45.3}$ $954^{+596}_{-546}\ (1.2\sigma)$ $103.2^{+191.0}_{-67.4}$ $49.2^{+26.9}_{-31.3}$
Velocity dispersion 100k $339^{+135}_{-127}\ (5.8\sigma)$ $240.9^{+47.9}_{-68.7}$ $54.5^{+22.7}_{-22.3}$ $1111^{+140}_{-135}\ (5.6\sigma)$ $258.0^{+13.5}_{-13.2}$ $48.8^{+8.1}_{-7.1}$
Scenario 1 (Pessimistic) $636^{+431}_{-388}\ (1.3\sigma)$ $208.6^{+92.4}_{-130.1}$ $53.1^{+25.2}_{-34.8}$ $1011^{+436}_{-416}\ (1.7\sigma)$ $256.9^{+68.4}_{-192.0}$ $60.2^{+20.4}_{-24.3}$
Scenario 2 (Realistic) $715^{+394}_{-356}\ (1.7\sigma)$ $214.4^{+51.7}_{-61.2}$ $44.1^{+28.5}_{-25.3}$ $1436^{+426}_{-422}\ (2.2\sigma)$ $262.8^{+22.9}_{-22.7}$ $36.3^{+17.5}_{-12.5}$
Scenario 3 (Ideal) $541^{+157}_{-156}\ (4.2\sigma)$ $237.5^{+67.9}_{-132.1}$ $69.4^{+14.1}_{-19.1}$ $1272^{+174}_{-176}\ (4.1\sigma)$ $257.8^{+20.8}_{-20.3}$ $57.4^{+10.9}_{-9.2}$
@@
For a fiducial value of $v_{o,\rm true}=1109.46$ $\mathrm{km}\,\mathrm{s}^{-1}$}, all three spectroscopic follow-up scenarios exclude the CMB-inferred value of $v_{o}$ with a significance ranging from $1.7\sigma$ in the pessimistic scenario to $4.1\sigma$ in the ideal scenario. Conversely, when the CMB-inferred value of $v_{o}$ is adopted as the fiducial velocity, the source number-count value is excluded with a significance ranging from $1.3\sigma$ to $4.2\sigma$ across the same scenarios. These results indicate that strong-lensing measurements could provide valuable complementary information on the cosmic dipole tension in realistic spectroscopic follow-up configurations, with a potential discrimination between the two values of $v_{o}$ at the $\sim 2\sigma$ level. In the presence of a dedicated spectroscopic follow-up of strong lenses, however, this probe could make a decisive contribution to this debate, with a discriminating power exceeding $4\sigma$.
@@ 5 Discussion
$\langle\theta^{\prime}_{E}\rangle$ $=\langle\theta_{E}\rangle\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)$ (5.2)
@@
$+\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}(\langle\theta_{E}\rangle-\theta^{\prime}_{E*})\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}f_{2}(z^{\prime}_{s},z^{\prime}_{l},\theta^{\prime}_{E*},F^{\prime}>F^{\prime}_{*})\mu(z^{\prime}_{s},z^{\prime}_{l},\theta^{\prime}_{E*},F^{\prime}_{*})\,.$
@@
The contribution from the size threshold, encoded in the third line of Eq. (5.2), is also suppressed, but this time by a factor $(\langle\theta_{E}\rangle-\theta^{\prime}_{E*})$. In this case, there is a balance between this factor and the parameter $\mu$. If most systems have sizes well above the threshold $\theta^{\prime}_{E*}$, then $(\langle\theta_{E}\rangle-\theta^{\prime}_{E*})$ will be large, but the effect will be strongly suppressed by $\mu$, which tends to zero in the limit where there are no rings below the threshold. Conversely, if most rings lie close to the threshold, then $\mu$ can be large, but the effect is suppressed by $(\langle\theta_{E}\rangle-\theta^{\prime}_{E*})$, which becomes small for rings near the threshold. We estimate the magnitude of this effect by evaluating $\mu$ from our sample, assuming a threshold of $\theta^{\prime}_{E*}=0.2\hbox{${}^{\prime\prime}$}$, and computing the third line of Eq. (5.2). We find that this contribution is of order $0.05\%$ of the main effect. Although significantly larger than the contribution from the flux threshold, it remains completely negligible compared to the main effect.
This demonstrates that the dipole in the Einstein rings is almost insensitive to the strong lensing selection function, provided it is uniform across the sky. As such, it provides a very robust way of measuring the observer velocity. In other words, neglecting these threshold effects would bias the measurement of the observer velocity by $0.05\%$ at most, largely below the statistical uncertainty.
@@
Finally, a main limitation of this study is that we ignore the impact of ellipticity in the mass distribution of the lensing galaxies and cosmic shear. The ellipticity of the Einstein ring contains additional directional information that could help constrain the dipole direction while simultaneously acting as a source of noise when combining lensing and kinematic information. In the lensing-only case, we emphasize that we use the azimuthally averaged Einstein radius as our primary observable. This quantity is sensitive to the enclosed mass of the galaxy, the angular-diameter–distance ratio, and cosmic convergence, which is expected to average out on large scales. In the lensing+kinematic case, ellipticity in the mass distribution of the lens galaxy introduces additional scatter in the predicted central velocity dispersion. Departures from isothermality in the mass profile would produce a similar effect. In this work, we account for both sources of uncertainty by adopting a $5\%$ modeling error in addition to the $5\%$ measurement error. Future improvements of this technique may include full dynamical modeling of the lenses as well as shear and convergence correction from weak lensing maps, allowing one to exploit the directional information encoded in the elliptical deformation of the Einstein ring. The aberration due to the observer’s velocity might appear as a dipole component in the line-of-sight shear of strong gravitational lenses, proposed by [Fleury2026].

### 19 Mar 2026 v2

@@ 3.3 Lensing only results
While the dipole is detected in both the “100k” and “400k” samples for the highest peculiar velocity, $v_{o,\rm true}=1109.46$$\mathrm{km}\,\mathrm{s}^{-1}$}, the lowest CMB-inferred value does not produce a sufficiently strong signal to be detected in the “100k” sample. Fixing the dipole direction leads to a moderate tightening of the constraints on the dipole amplitude, improving the 68% credible interval width on $v_{o}$ by approximately 10 to 30%, depending on the sample size and fiducial velocity.
We therefore find that, when relying on lensing information alone, the Euclid lens sample is unlikely to provide strong discriminatory power between the CMB-inferred dipole amplitude and the larger values suggested by source number count analyzes. The resulting constraints are not sufficiently precise to exclude the CMB-inferred value beyond the $\sim 1.2\sigma$ level, even when fixing the dipole direction. For the “400k” sample, the discriminating power reaches $\sim 2\sigma$. We explore the possibility of including the stellar kinematic information to improve the precision of this measurement in the next section.
@@
Lensing only 100k $967^{+1065}_{-674}$ $273.7^{+56.3}_{-211.8}$ $14.7^{+42.3}_{-39.9}$ $1527^{+1157}_{-949}$ $287.8^{+39.8}_{-151.4}$ $29.4^{+34.5}_{-25.9}$
Lensing only 400k $347^{+419}_{-243}$ $238.2^{+84.7}_{-178.9}$ $12.0^{+47.3}_{-50.7}$ $900^{+511}_{-454}$ $278.0^{+42.3}_{-101.8}$ $43.2^{+29.3}_{-24.8}$
Lensing only 100k - Fixed direction $940^{+694}_{-582}$ $1541^{+787}_{-760}$
Lensing only 400k - Fixed direction $391^{+344}_{-254}$ $1002^{+391}_{-396}$
@@ 4.1 Spectroscopic velocity dispersion
$\theta^{\prime}_{E,\rm spec}=\frac{4\pi\sigma_{v}^{2}}{c^{2}}\frac{d[z_{l}^{\prime},z_{s}^{\prime}]}{d[z_{s}^{\prime}]}\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}+\frac{D_{LS}}{d[z_{l}^{\prime},z_{s}^{\prime}]}\frac{v_{o}}{c}-\frac{D_{S}}{d[z_{s}^{\prime}]}\frac{v_{o}}{c}\right),$ (4.11)
@@ 4.2 Fundamental Plane
$\log_{10}R_{e}$ $=a\log_{10}\sigma_{\mathrm{FP}}+b\log_{10}I_{e}+c\,.$ (4.13)
Only the angular size of the effective radius can be directly measured from the imaging data, but $R_{e}$ can be straightforwardly derived from this measurement and from the (photometric) redshift of the lens, assuming a cosmological model. We assume $a=1.177$, b=0.793b=-0.793 and $c=0.5$, obtained from the latest DESI measurement [Said2025]. Measurement errors on these parameters are ignored, as they lead to errors much smaller than the intrinsic scatter of the FP. While $I_{e}$ is invariant under a boost (as shown in the App. D), $R_{e}$ is not. It can be computed from the observed subtended angle $\alpha^{\prime}$ and lens redshift $z_{l}^{\prime}$, which are affected by the boost, as
$R_{e}=\alpha d_{l}=\alpha^{\prime}d[z_{l}^{\prime}]\left(1+\frac{v_{o}}{c}\cos(\theta_{\rm cm}^{\prime})+\frac{D_{L}}{d[z_{l}^{\prime}]}\frac{v_{o}}{c}\right)\,.$ (4.14)
@@
$\theta^{\prime}_{E,\rm FP}=\frac{4\pi\sigma_{\rm FP}^{\prime 2}}{c^{2}}\frac{d[z_{l}^{\prime},z_{s}^{\prime}]}{d[z_{s}^{\prime}]}\left(1-\frac{v_{o}}{c}\cos(\theta^{\prime}_{\rm cm})+\frac{D_{LS}}{d[z_{l}^{\prime},z_{s}^{\prime}]}\frac{v_{o}}{c}-\frac{D_{S}}{d[z_{s}^{\prime}]}\frac{v_{o}}{c}+2\Sigma_{v}\frac{v_{o}}{c}\right)\,$, (4.15)
@@
$\Sigma_{v}$ $=\left(\frac{1}{a}\cos(\theta_{\rm cm}^{\prime})+\frac{D_{L}}{ad[z_{l}^{\prime}]}\right)\,.$ (4.17)
@@
Details on how the uncertainties $\delta\sigma_{\mathrm{FP}}$ are calculated can be found in App. C. Similarly to Eq. (4.12), we introduced a bias and systematic error term to account for model mismatch, peculiar velocity effects, and redshift measurement errors when converting Einstein angles to velocity dispersions. We treat separately the likelihood contributions from lenses with only photometric redshift estimates and those with spectroscopic redshift measurements for both source and lens, since these two subsamples require different bias and systematic error corrections due to the differing precision of their redshift determinations.
@@ 4.3 Euclid Forecast
Finally, we assess the constraining power of several spectroscopic follow-up scenarios for the Euclid strong-lens sample. In a first, pessimistic scenario, we assume that only 15 00015\,000 lenses receive spectroscopic follow-up, either from large spectroscopic surveys such as DESI [DESI2016] or 4MOST [4MOSToverview, 4SLSLS], or from dedicated follow-up programs. While both redshift and stellar velocity-dispersion measurements are, in principle, accessible with fiber spectroscopy, the latter are significantly more challenging, as they require a higher signal-to-noise ratio to robustly detect the absorption features of the lens galaxy. As a result, even if a lens is targeted by DESI or 4MOST, a substantial fraction of systems is expected to yield redshift measurements only. In this pessimistic scenario, we therefore assume that velocity-dispersion measurements are successfully obtained for only a third of the spectroscopically targeted lenses, i.e. 5 0005\,000 lenses. The remaining lenses are assumed to have spectroscopic, photometric, or no redshift information, as summarized in Table 3.
@@
Lensing only 100k $966^{+1094}_{-685}\ (0.8\sigma)$ $276.3^{+53.8}_{-208.1}$ $13.9^{+41.9}_{-39.8}$ $1311^{+1198}_{-875}\ (1.1\sigma)$ $44.6^{+46.6}_{-73.3} 22$$2.7^{+35.9}_{-30.1}$
FP (Photo-z) 100k $1132^{+662}_{-621}\ (1.4\sigma)$ $229.7^{+33.5}_{-38.0}$ $26.0^{+28.0}_{-18.7}$ $1347^{+685}_{-735}\ (1.3\sigma)$ $216.6^{+27.1}_{-27.4}$ $11.7^{+15.9}_{-20.3}$
FP (Spectro-z) 100k $444^{+543}_{-316}\ (1.6\sigma)$ $171.9^{+75.4}_{-72.8}$ $6.4^{+45.3}_{-44.7}$ $828^{+533}_{-474}\ (1.8\sigma)$ $266.8^{+48.1}_{-106.7}$ $39.5^{+31.6}_{-27.1}$
Velocity dispersion 100k $340^{+120}_{-110}\ (7.2\sigma)$ $256.3^{+31.2}_{-36.5}$ $45.1^{+22.7}_{-16.8}$ $1116^{+109}_{-103}\ (7.3\sigma)$ $258.6^{+10.5}_{-10.5}$ $48.7^{+6.0}_{-5.5}$
Scenario 1 (Pessimistic) $888^{+482}_{-451}\ (1.8\sigma)$ $270.5^{+32.4}_{-40.4}$ $28.6^{+27.5}_{-19.2}$ $809^{+491}_{-444}\ (2.1\sigma)$ $306.9^{+31.5}_{-135.4}$ $29.2^{+30.6}_{-20.4}$
Scenario 2 (Realistic) $614^{+256}_{-252}\ (2.8\sigma)$ $222.0^{+63.0}_{-87.3}$ $58.8^{+21.0}_{-24.7}$ $1154^{+293}_{-277}\ (2.8\sigma)$ $225.7^{+29.1}_{-29.3}$ $51.3^{+18.0}_{-14.3}$
Scenario 3 (Ideal) $506^{+119}_{-121}\ (5.5\sigma)$ $238.7^{+57.4}_{-93.7}$ $70.1^{+13.6}_{-17.3}$ $1213^{+143}_{-136}\ (5.3\sigma)$ $253.6^{+15.0}_{-14.9}$ $54.2^{+8.1}_{-7.4}$
@@
For a fiducial value of $v_{o,\rm true}=1109.46$ $\mathrm{km}\,\mathrm{s}^{-1}$}, all three spectroscopic follow-up scenarios exclude the CMB-inferred value of $v_{o}$ with a significance ranging from $2.1\sigma$ in the pessimistic scenario to $5.3\sigma$ in the ideal scenario. Conversely, when the CMB-inferred value of $v_{o}$ is adopted as the fiducial velocity, the source number-count value is excluded with a significance ranging from $1.8\sigma$ to $5.5\sigma$ across the same scenarios. These results indicate that strong-lensing measurements could provide valuable complementary information on the cosmic dipole tension, with a potential discrimination between the two values of $v_{o}$ at the $\sim 2.8\sigma$ level, in a realistic spectroscopic follow-up configuration. In the presence of a dedicated spectroscopic follow-up of strong lenses, however, this probe could make a decisive contribution to this debate, with a discriminating power exceeding $5\sigma$.
@@ 5 Discussion
$\langle\theta^{\prime}_{E}\rangle$ $=\langle\theta_{E}\rangle\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$ (5.2)
@@
$+\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}(\langle\theta_{E}\rangle-\theta^{\prime}_{E*})\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}f_{2}(z^{\prime}_{s},z^{\prime}_{l},\theta^{\prime}_{E*},F^{\prime}>F^{\prime}_{*})\mu(z^{\prime}_{s},z^{\prime}_{l},\theta^{\prime}_{E*},F^{\prime}_{*})\,.$
@@
The contribution from the size threshold, encoded in the third line of Eq. (5.2), is also suppressed, but this time by a factor $(\langle\theta_{E}\rangle-\theta^{\prime}_{E*})$. In this case, there is a balance between this factor and the parameter $\mu$. If most systems have sizes well above the threshold $\theta^{\prime}_{E*}$, then $(\langle\theta_{E}\rangle-\theta^{\prime}_{E*})$ will be large, but the effect will be strongly suppressed by $\mu$, which tends to zero in the limit where there are no rings below the threshold. Conversely, if most rings lie close to the threshold, then $\mu$ can be large, but the effect is suppressed by $(\langle\theta_{E}\rangle-\theta^{\prime}_{E*})$, which becomes small for rings near the threshold. We estimate the magnitude of this effect by evaluating $\mu$ from our sample, assuming a threshold of $\theta^{\prime}_{E*}=0.2\hbox{${}^{\prime\prime}$}$, and computing the third line of Eq. (5.2). We find that this contribution is of order $0.1\%$ of the main effect. Although significantly larger than the contribution from the flux threshold, it remains completely negligible compared to the main effect.
This demonstrates that the dipole in the Einstein rings is almost insensitive to the strong lensing selection function, provided it is uniform across the sky. As such, it provides a very robust way of measuring the observer velocity. In other words, neglecting these threshold effects would bias the measurement of the observer velocity by $0.1\%$ at most, largely below the statistical uncertainty.
@@
Finally, a limitation of this study is that we ignore the impact of ellipticity in the mass distribution of the lensing galaxies. However, we emphasize that we use the azimuthally averaged Einstein radius as our primary observable. This quantity is sensitive to the enclosed mass of the galaxy, the angular-diameter distance ratio, and cosmic convergence, which is expected to average out on large scales but is still included as an additional noise contribution to the Einstein radius measurement. In the lensing+kinematic case, ellipticity in the mass distribution of the lens galaxy introduces additional scatter in the predicted central velocity dispersion. Departures from isothermality in the mass profile would produce a similar effect. In this work, we account for both sources of uncertainty by adopting a $5\%$ modeling error in addition to the $5\%$ measurement error. Future analyzes on real data may include full dynamical modeling of the lenses, as well as shear and convergence corrections from weak lensing maps.